# Bell Nonlocality from Wigner Negativity in Qudit Stabilizer States

### v3

@@ Abstract
with a function $f(k)$ and derive an analytic expression that map Pauli to Clifford operators, where $f(k)$ turns out to be a polynomial of degree $3$. In the following, we refer to their generalized $T$-gates as unitary cube operators.
For $d>3$, it is $f(k)=\nu_{k}=12^{-1}k(\gamma+k(6z+(2k+3)\gamma))+\epsilon k$ with $z,\epsilon\in\mathbb{Z}_{d}$ and $\gamma\in\mathbb{Z}_{d}^{\ast}=\mathbb{Z}_{d}\setminus\{0\}$. We discuss the case $d=3$ later together with non-character unitary operators. The coefficients are such that $U^{\phantom{\dagger}}_{\nu}XU^{\dagger}_{\nu}=\omega^{\epsilon}XZ^{z}\sum_{k\in\mathbb{Z}_{d}}\omega^{2^{-1}\gamma k^{2}}|k\rangle\langle k|$.
@@
Stabilizer states transformed by the adjoint action of unitary cube operators can then showcase Wigner negativity. A measure of the amount of Wigner negativity is its volume defined as minus the sum of all negative values, or, since the Wigner function is normalized, as $N[\rho]=(\sum_{u}|W_{u}(\rho)|-1)/2\,.$ For the Bell state $|\Phi\rangle=\sum_{k=0}^{d-1}|k\,k\rangle/\sqrt{d}$, this leads to $|\Phi_{\nu}\rangle:=\left(U_{\nu}\otimes\mathds{1}\right)|\Phi\rangle\,,$ with the Wigner function, for $d>3$,
$W_{u_{1},u_{2}}(|\Phi_{\nu}\rangle\langle\Phi_{\nu}|)=\frac{1}{d^{3}}\delta_{(u_{1})_{x},(u_{2})_{x}}\sum^{d-1}_{k=0}\omega^{a_{3}k^{3}+a_{1}k}\,.$ (7)
@@
Note that $a_{1}(u_{1},u_{2})=\epsilon+(u_{1})_{z}+(u_{2})_{z}+z(u_{1})_{x}+2^{-1}\gamma((u_{1})_{x}^{2}-(u_{1})_{x}+6^{-1})$ and $a_{3}=24^{-1}\gamma$, such that, independently of $u_{1}$ and $u_{2}$, $a_{1}$ can take any value in $\mathbb{Z}_{d}$ and $a_{3}$ any value in $\mathbb{Z}_{d}^{\ast}$. For $(u_{1})_{x}=(u_{2})_{x}$, the value of $W^{\nu}_{u_{1},u_{2}}:=W_{u_{1},u_{2}}(|\Phi_{\nu}\rangle\langle\Phi_{\nu}|)$ is determined by the sum of characters with a third-order odd polynomial as an argument. Although such polynomials are difficult to analyze for general odd prime $d$, a simple brute-force search finds values $a_{1},a_{3}$ such that $W^{\nu}_{u_{1},u_{2}}<0$ for any $u_{1}$ and $u_{2}$. The negativity of $W^{\nu}_{u_{1},u_{2}}$ is pivotal in achieving a Bell violation with the Bell operator in Eq. (11).
$|W^{\nu}_{u_{1},u_{2}}|\leq\frac{2}{d^{2}\sqrt{d}}\,.$ (8)
@@
Table 1: Extremal values of the Wigner function $W^{\nu}_{v_{1},v_{2}}$ in Eq. (7) and Eq. (21), maximal values of its negativity volume $N^{\nu}:=N[\Phi_{\nu}]$ and the nonlocal bound $B^{\max}_{\mathrm{lhv}}$ from Eq. (14), all found through an exhaustive search, as well as values of $C_{\min}$ (Eq. (9)) for small prime dimension $d$. For $d\leq 7$, we saturate the bound of Eq. (10) and achieve $C_{\min}=d^{3}\min W^{\nu}_{v_{1},v_{2}}$. * For $d=3$, we instead use $\chi(f)=1+2\cos\left(8\pi/9\right)$ as we allow for cube roots of the character. ** For $d=2$, we use the Wigner function adapted to qubits for the Bell state and the $T$-gate in Eq. (28). Its minimum value is $-1/\sqrt{2}\approx-0.707$, corresponding to Tsirelson’s bound Cirelson1980 . Anyway, the Bell operator (11) and analysis thereof is not apt to qubits.
$\min_{g}\chi(g)\geq C_{\min}=1-(d-1)\cos\left(\pi/d\right)$ (9)
Bell inequality.
We propose a novel Bell operator in which a specific state $\sigma$ becomes first a witness for contextuality and Wigner negativity and ,then, for nonlocality. The operator, corresponding to a entangled two-qudit state $\sigma$ is
$\mathcal{B}[\sigma]=\sigma=\frac{1}{d^{2}}\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}\chi_{u_{1},u_{2}}(\sigma)\,T_{u_{1}}\otimes T_{u_{2}}\,,$ (11)
with the state’s characteristic function $\chi_{u_{1},u_{2}}(\sigma)=\mathrm{Tr}(\sigma\,T_{-u_{1}}\otimes T_{-u_{2}})$.
The corresponding measurements are defined in terms of projectors $\Pi^{(v)}_{m}=d^{-1}\sum_{k\in\mathbb{Z}_{d}}\omega^{mk}T_{v}^{k}$ of outcome $m\in\mathbb{Z}_{d}$ and $v\in\mathbb{Z}_{d}^{2}$ on the eigenbasis of the Pauli operators. Although the operator comprises of $d^{4}$ local operators $T_{v}$, only $(d+1)^{2}$ measurements corresponding to mutually unbiased bases for each qudit are necessary.
$\langle\mathcal{B}[\sigma]\rangle_{\mathrm{nc}^{\ast}}$ $=\frac{1}{d^{2}}\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}\chi_{u_{1},u_{2}}(\sigma)\,\omega^{[a_{1},{u_{1}}]}\omega^{[a_{2},{u_{2}}]}$
$=d^{2}\,W_{a_{1},a_{2}}(\sigma)\,.$ (12)
$\left\langle\mathcal{B}[\sigma]\right\rangle_{\mathrm{nc}}\leq d^{2}\max_{a_{1},a_{2}\in\mathbb{Z}_{d}^{2}}|W_{a_{1},a_{2}}(\sigma)|\equiv B_{\mathrm{nc}}^{\max}\,.$ (13)
On the other hand, a pure state state $\sigma$ maximally violates the non-contextual inequality, which it defines via its characteristic function, among all qudit states if $\mathrm{Tr}(\sigma^{2})=1>B_{\mathrm{nc}}^{\max}$. We find with Eq. (31) that non-trivial negativity volume, $N[\Psi]>0$, is necessary for such a violation. As a result, Eq. (11) defines a family of inequalities based on Wigner negative states that are witnesses of contextuality if $1>d^{2}\max_{a_{1},a_{2}\in\mathbb{Z}_{d}^{2}}|W_{a_{1},a_{2}}[\sigma]|$. This specifically exploit the equivalence of Wigner negativity and contextuality with Pauli measurements Delfosse2017 .
$\langle\mathcal{B}[\sigma]\rangle_{\mathrm{lhv}}\leq\max_{\alpha,\beta\in\mathbb{Z}_{d}^{d+1}}\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}\chi_{u_{1},u_{2}}[\sigma]\,\omega^{\alpha_{r(u_{1})}+\beta_{r(u_{2})}}\equiv B^{\max}_{\mathrm{lhv}}\,.$ (14)
In terms of the Wigner function, the Bell bound is
$B^{\max}_{\mathrm{lhv}}=\max_{\alpha,\beta\in\mathbb{Z}_{d}^{d+1}}\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}W_{u_{1},u_{2}}[\sigma]F(\alpha)_{u_{1}}F(\beta)_{u_{1}}\,.$ (15)
with the discrete Fourier transform of $\omega^{a_{r(u)}}$, $F(a)_{u}=\sum_{v\in\mathbb{Z}_{d}^{2}}\omega^{[u,v]+\alpha_{r(v)}}/d$. The values of $F(\alpha)_{u}$ range between $-1,0,\dots,d$, the sum of entries is $\sum_{u\in\mathbb{Z}_{d}^{2}}F(\alpha)_{u}=d$, and the vector’s norm is equally $\|F(\alpha)\|_{2}=d$. For an $\alpha\in\mathbb{Z}_{d}^{d+1}$ such that $F(\alpha)_{v}=d$ for some $v\in\mathbb{Z}_{d}^{2}$, it is $\alpha_{r(u)}=[a,u]$ for some $a\in\mathbb{Z}_{d}^{2}$, and $F(\alpha)_{w}=0$ for all $w\in\mathbb{Z}_{d}^{2}\setminus\{v\}$, which is precisely, the non-contextual value assignment used for the Wigner negativity witness or non-contextual inequality (13). To evaluate the Bell bound (15) further, the Cauchy-Schwarz inequality gives $B^{\max}_{\mathrm{lhv}}[\sigma]\leq\|W[\sigma]\|_{2}\|F(\alpha)\|_{2}\|F(\beta)\|_{2}\leq d^{2}\|W[\sigma]\|_{2}$. To observe a Bell violation, it is necessary that $B_{\mathrm{lhv}}[\sigma]<1$, which requires at least $\|W[\sigma]\|_{2}\leq 1$. Since the non-contextual value assignments are a special case of the deterministic lhv models’, it is $B^{\max}_{\mathrm{nc}}\leq B^{\max}_{\mathrm{lhv}}$ such that Wigner negativity of the pure resource state $\sigma$ as well as the state used for the quantum realization is again essential for nonlocality. Whether the state defines a notable Bell inequality beyond these criteria requires an analysis of the state’s characteristic or Wigner function. Bell states rotated by local unitary cube operators $|\Phi_{\nu}\rangle$, also employed in Howard2015bell , are prime candidates to define a Bell inequality in terms of their characteristic function and Bell bounds in Tab. 1.
Qudit stabilizer states, however, do not exhibit Wigner negativity. To construct a Bell inequality for qudit stabilizer states, we measure the operators that make up the Wigner function $W^{\nu}_{v_{1},v_{2}}$ to exploit the Wigner negativity of the stabilizer states under unitary cube operators. To highlight the negative values, a favorable coefficient distinguishes the negative values from the positive ones, the head-on candidate being the Wigner function itself. For the qudit Bell state the resulting operator is $\mathcal{B}_{\nu}=\sum_{v_{1},v_{2}\in\mathbb{Z}_{d}^{2}}W^{\nu}_{v_{1},v_{2}}\phantom{\cdot}U^{\phantom{\dagger}}_{\nu}A_{v_{1}}U^{\dagger}_{\nu}\otimes A_{v_{2}}$. In terms of qudit Pauli operators, it is
$\mathcal{B}_{\nu}=\frac{1}{d^{2}}\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}\chi^{\nu}_{u_{1},u_{2}}\,U^{\phantom{\dagger}}_{\nu}T_{u_{1}}U^{\dagger}_{\nu}\otimes T_{u_{2}}\,,$ (16)
with the characteristic function $\chi^{\nu}_{u_{1},u_{2}}=\langle\Phi_{\nu}|T_{-u_{1}}\otimes T_{-u_{2}})|\Phi_{\nu}\rangle$. The non-contextual bounds arise for the same reasons as those underlying the inequality (13) such that $\left\langle\mathcal{B}_{\nu}\right\rangle_{\mathrm{nc}}\leq d^{2}\max_{v_{1},v_{2}\in\mathbb{Z}_{d}^{2}}|W^{\nu}_{v_{1},v_{2}}|\equiv B_{\mathrm{nc}}^{\max}\,.$ The Bell operator’s expectation value for the Bell state is $\mathrm{Tr}\left(\mathcal{B}_{\nu}|\Phi\rangle\langle\Phi|\right)=\sum_{v_{1},v_{2}\in\mathbb{Z}_{d}^{2}}d^{2}(W^{\nu}_{v_{1},v_{2}})^{2}=1\,,$ since the rotated Bell state is a pure state. Moreover, we show in Eq. (26) that the expectation value of the Bell operator for any bipartite state $\sigma$ is a measure for the singlet fraction
$\langle\mathcal{B}_{\nu}\rangle_{\sigma}=\mathrm{Tr}\left(\mathcal{B}_{\nu}\sigma\right)=\langle\Phi|\sigma|\Phi\rangle\,.$ (17)
From Eq. (32), the expectation value with a quantum state $\sigma$ can be bound by $B_{\mathrm{nc}}^{\max}$ and the volume of Wigner negativity using Hölder’s inequality with the 1-norm and the maximum norm,
$\mathrm{Tr}\left(\mathcal{B}_{\nu}\sigma\right)\leq B_{\mathrm{nc}}^{\max}(1+2N[\sigma_{\nu}])\,,$ (18)
with $\sigma_{\nu}:=(U_{\nu}\otimes\mathds{1})^{\dagger}\sigma(U_{\nu}\otimes\mathds{1})$. In particular, it is $1\leq B_{\mathrm{nc}}^{\max}(1+2N[\Phi_{\nu}])\,.$ Thus, a non-contextual violation $\mathrm{Tr}\left(\mathcal{B}_{\nu}|\Phi\rangle\langle\Phi|\right)=1>B_{\mathrm{lhv}}^{\max}$ is possible for a nontrivial negativity volume, $N[\Phi_{\nu}]>0$. The previous analysis of character polynomials defining the Wigner function $W^{\nu}_{v_{1},v_{2}}$, in particular Eq. (8), gives
$B_{\mathrm{nc}}^{\max}\leq 2/\sqrt{d}$ (19)
for $d>3$. For $d=3$, an exhaustive search using Eq. (21) finds $B_{\mathrm{nc}}^{\max}<0.844$. Table 1 lists such values $d^{2}\max W^{\nu}_{v_{1},v_{2}}[\Phi_{\nu}]$ for $d\leq 23$. For a classical model, the correlations from $\mathcal{B}_{\nu}$ are equivalent to those from $\mathcal{B}[\Phi_{\nu}]$. The maximum value attainable by an lhv model is the same and an exhaustive search finds tight values for $B^{\max}_{\mathrm{lhv}}$ for small odd prime $d$ in Tab. 1, which seem to decrease with growing $d$.
For a more compact Bell operator, we now focus solely on the stabilizer elements, $T_{(x,z)}\otimes T_{(x,-z)}|\Phi\rangle=|\Phi\rangle$, which reduces the number of measurements from $(d+1)^{2}$ to $d^{2}+1$, and leads to
$\mathcal{B}_{\nu}^{\prime}=\sum_{x,z,t\in\mathbb{Z}_{d}}W^{\nu}_{(x,z),(x,-z)}\phantom{\cdot}U_{\nu}A_{(x,z)}U^{\dagger}_{\nu}\otimes A_{(x,t-z)}\,.$ (20)
For convenience, we use the operator $A_{u}$, however, the measurements are still corresponding to the qudit Pauli operators $T_{u}$. The expectation values have the same form as for the full Bell operator (16), but the summation and the maximum take only coefficients $x,z,t\in\mathbb{Z}_{d}$. The Bell operator is a measure of the singlet fraction and the volume of Wigner negativity. The Bell inequality is structurally the same and Bell state achieves the same Bell violation.
As a result, we construct a family of Bell inequalities based on the Wigner function of the Bell state rotated by cube unitary operators. In Appendix B, we outline how these can be regarded as a natural generalization of the Clauser-Horne-Shimony-Holt (CHSH) inequality.
@@
where $X_{(1/3)}=UXU^{\dagger}$, and $U=\mathrm{diag}(1,\omega^{2/3},\omega^{1/3})$. Note that these operators $O=X,X_{(1/3)}$ are not Hermitian but unitary yet represent valid measurements in terms of their Fourier transform with projectors $\Pi_{m}=d^{-1}\sum_{k\in\mathbb{Z}_{d}}\omega^{mk}O^{k}$ of outcome $m\in\mathbb{Z}_{d}$. The same holds for a deterministic value assignment. As a result, any lhv model attains $B_{3,\mathrm{lhv}}^{\max}=\max(\omega^{a_{0}+b_{0}}+\omega^{2a_{0}+2b_{1}}+\omega^{2a_{1}+2b_{0}}+\omega^{a_{1}+b_{1}+1}+\,h.c.)=6+\omega+\omega^{2}=5$, where the maximum is taken over $a_{0},b_{0},a_{1},b_{1}\in\mathbb{Z}_{3}$, and is attained for $b_{1}=2$ and $a_{0},b_{0},a_{1}=0$. On the other hand, the quantum expectation value, $\mathrm{Tr}\left(\mathcal{B}_{3}\,|\Phi\rangle\langle\Phi|\right)=1+3(2\omega^{1/3}+\omega^{-2/3})+\,h.c.\,\approx 5.412\,,$ exceeds that of the classical model, $B_{3,\mathrm{lhv}}^{\max}<\mathrm{Tr}\left(\mathcal{B}_{3}\,|\Phi\rangle\langle\Phi|\right)$.
@@
Multipartite Bell inequality.
Lastly, we generalize the Bell operator in Eq. (20) to any multipartite qudit stabilizer state. Consider the state $|S\rangle$ in $n$ qudits with stabilizer elements $S_{\bm{u}}=\omega^{[{\bm{a}},{\bm{u}}]}T_{\bm{u}}$ for all ${\bm{u}}\in\Sigma\subset\mathbb{Z}_{d}^{2n}$. For a unitary cube operator $U_{\nu}$ acting on the first qudit,
$\mathcal{B}_{S}=\sum_{{\bm{u}}\in\Sigma,t\in\mathbb{Z}_{d}}W^{\nu}_{{\bm{u}}}\left(U^{\phantom{\dagger}}_{\nu}\,A_{u_{1}+(0,t)}\,U^{\dagger}_{\nu}\right)\bigotimes_{i=2}^{n}A_{u_{i}}\,.$ (24)
@@
We have constructed several families of qudit Bell inequalities for any odd prime dimension. First, we propose a novel form of general Bell inequalities where a bipartite entangled state exhibiting Wigner negativity acts as a witness for nonlocality. Then, we construct a family of Bell operators aiming at a nonlocal violation with the Bell state. Similarly to the $\pi/8$-rotation leading to the CHSH inequality, our Bell operator consists of stabilizer elements under local rotation of a generalization of the qubit $T$-gate characterized in Howard2012 . In Eq. (29), we illustrate how the Bell inequalities can be regarded as a natural extension of the CHSH inequality. The constituting measurements test correlations related to the rotated states’ Wigner negativity, a witness of non-classicality. The maximally entangled Bell state violates the corresponding Bell inequality and attains the quantum bound due to the separation between the 1-norm and the maximum norm of its rotated Wigner function. Furthermore, the Bell operator measures the singlet fraction and bounds the extent of Wigner negativity. Our family of Bell operators, as well as its quantum and classical bound, is easy to analyze and can be adapted to any entangled stabilizer state, which also reduces the number of measurements. The measurements can be generalized to any non-Clifford unitary operator, and we specifically discuss those from higher-degree polynomials over finite fields, which we find to achieve larger Wigner negativity and thus increase the separation between quantum and classical models. Moreover, nonlocal correlations can emerge with qudit Pauli operators when rotated by unitary operators with a non-character spectrum in bipartite systems. These operators have an obvious advantage, which is resorting to phases $\omega^{q}$ that are beyond the description of any local value assignment that can only resort to characters $\omega^{k}$. Such operations might arise from interactions with higher-dimensional systems, but the feasibility of the implementation of any non-Clifford operator present in this work remains an open question. Its study heavily depends on the platform of experimental realization with possible candidates being super-conducting circuits, cold atoms and optical systems involving GKP (Gottesman–Kitaev–Preskill) states PhysRevA.64.012310 .

### v4

@@ Abstract
with a function $f(k)$ and derive an analytic expression that map Pauli to Clifford operators, where $f(k)$ turns out to be a polynomial of degree $3$ for $d>3$. We discuss the case $d=3$ later together with non-character unitary operators. Qudit Clifford operators were characterized by Appleby appleby2005symmetric , Appleby2009 who finds a unique representation for every Clifford operator with a decomposition into a Pauli operator and a unitary evolution under quadratic Hamiltonians. Specifically, it is $f(k)=\nu_{k}=12^{-1}k(\gamma+k(6z+(2k+3)\gamma))+\epsilon k$ with $z,\epsilon\in\mathbb{Z}_{d}$ and $\gamma\in\mathbb{Z}_{d}^{\ast}=\mathbb{Z}_{d}\setminus\{0\}$ for $d>3$. The coefficients are such that $U^{\phantom{\dagger}}_{\nu}XU^{\dagger}_{\nu}=\omega^{\epsilon}XZ^{z}\sum_{k\in\mathbb{Z}_{d}}\omega^{2^{-1}\gamma k^{2}}|k\rangle\langle k|$. In the following, we refer to the generalized $\pi/8$-gates as unitary cube operators.
@@
Stabilizer states transformed by the adjoint action of unitary cube operators can then showcase Wigner negativity. A measure of the amount of Wigner negativity is its volume defined as minus the sum of all negative values, or, since the Wigner function is normalized, as $N[\rho]=(\sum_{u}|W_{u}[\rho]|-1)/2\,.$ For the Bell state $|\Phi\rangle=\sum_{k=0}^{d-1}|k\,k\rangle/\sqrt{d}$, this leads to $|\Phi_{\nu}\rangle:=\left(U_{\nu}\otimes\mathds{1}\right)|\Phi\rangle\,,$ with the Wigner function $W^{\nu}_{u_{1},u_{2}}:=W_{u_{1},u_{2}}[|\Phi_{\nu}\rangle\langle\Phi_{\nu}|]$,
$W^{\nu}_{u_{1},u_{2}}=\frac{1}{d^{3}}\delta_{(u_{1})_{x},(u_{2})_{x}}\sum^{d-1}_{k=0}\omega^{a_{3}k^{3}+a_{1}k}\,,$ (18)
@@
for $d>3$. Note that $a_{1}(u_{1},u_{2})=\epsilon+(u_{1})_{z}+(u_{2})_{z}+z(u_{1})_{x}+2^{-1}\gamma((u_{1})_{x}^{2}-(u_{1})_{x}+6^{-1})$ and $a_{3}=24^{-1}\gamma$, such that, independently of $u_{1}$ and $u_{2}$, $a_{1}$ can take any value in $\mathbb{Z}_{d}$ and $a_{3}$ any value in $\mathbb{Z}_{d}^{\ast}$. For $(u_{1})_{x}=(u_{2})_{x}$, the value of $W^{\nu}_{u_{1},u_{2}}$ is determined by the sum of characters with a third-order odd polynomial as an argument. Although such polynomials are difficult to analyze for general odd prime $d$, a simple brute-force search finds values $a_{1},a_{3}$ such that $W^{\nu}_{u_{1},u_{2}}<0$ for any $u_{1}$ and $u_{2}$. The negativity of $W^{\nu}_{u_{1},u_{2}}$ is pivotal in achieving a Bell violation with the Bell operator in Eq. (9). In Appendix A, we discuss further properties of this Wigner function and generalizations to achieve more negativity. For instance, the analysis of character polynomials defining the Wigner function $W^{\nu}_{v_{1},v_{2}}$, in particular Eq. (25), with $U=U_{\nu}$ gives
$B_{\mathrm{nc}}^{\max}\leq 2/\sqrt{d}$ (19)
for $d>3$. For $d=3$, an exhaustive search using Eq. (22) finds $B_{\mathrm{nc}}^{\max}<0.844$. Table 1 lists such values $d^{2}\max_{v_{1},v_{2}\in\mathbb{Z}_{d}^{2}}W^{\nu}_{v_{1},v_{2}}$ for $d\leq 23$. For a classical model, the correlations from $\mathcal{B}_{\nu}$ are equivalent to those from $\mathcal{B}[\Phi_{\nu}]$. The maximum value attainable by an lhv model is the same and an exhaustive search finds tight values for $B^{\max}_{\mathrm{lhv}}$ for small odd prime $d$ in Tab. 1, which seem to decrease with growing $d$.
For a more compact Bell operator, we solely consider the stabilizer elements, $T_{(x,z)}\otimes T_{(x,-z)}|\Phi\rangle=|\Phi\rangle$, which reduces the number of measurements from $(d+1)^{2}$ to $d^{2}+1$, and leads to
$\mathcal{B}_{\nu}^{\prime}=\sum_{x,z,t\in\mathbb{Z}_{d}}\chi^{\nu}_{(x,z),(x,-z)}\phantom{\cdot}U_{\nu}T_{(x,z)}U^{\dagger}_{\nu}\otimes T_{(x,t-z)}\,.$ (20)
The expectation values have the same form as for the full Bell operator (14), but the summation and the maximum take only coefficients $x,z,t\in\mathbb{Z}_{d}$. The Bell operator is a measure of the singlet fraction and the volume of Wigner negativity. The Bell inequality is structurally the same and Bell state achieves the same Bell violation. In Appendix B, we outline how these can be regarded as a natural generalization of the Clauser-Horne-Shimony-Holt (CHSH) inequality.
@@
Table 1: Extremal values of the Wigner function $W^{\nu}_{v_{1},v_{2}}$ in Eq. (18) and Eq. (22), maximal values of its negativity volume $N^{\nu}:=N[\Phi_{\nu}]$ and the nonlocal bound $B^{\max}_{\mathrm{lhv}}$ from Eq. (12), all found through an exhaustive search, as well as values of $C_{\min}$ (Eq. (26)) for small prime dimension $d$. For $d\leq 7$, we saturate the bound of Eq. (27) and achieve $C_{\min}=d^{3}\min W^{\nu}_{v_{1},v_{2}}$. * For $d=3$, we instead use $\chi(f)=1+2\cos\left(8\pi/9\right)$ as we allow for cube roots of the character. ** For $d=2$, we use the Wigner function adapted to qubits for the Bell state and the $\pi/8$-gate in Eq. (33). Its minimum value is $-1/\sqrt{2}\approx-0.707$, corresponding to Tsirelson’s bound Cirelson1980 . Anyway, the Bell operator (9) and analysis thereof is not apt to qubits.
For a general stabilizer state, we define the set ${\bm{u}}\in\Sigma\subset\mathbb{Z}_{d}^{2n}$ so that $\omega^{t({\bm{u}})}T_{\bm{u}}\in\mathcal{S}$ for a linear function $t$, which can be written as $t({\bm{u}})=[{\bm{a}},{\bm{u}}]$ for an ${\bm{a}}\in\mathbb{Z}_{d}^{2n}$. The stabilizer elements act trivially on the stabilizer state $\omega^{t({\bm{u}})}T_{{\bm{u}}}|S\rangle=|S\rangle$, while for any other $\omega^{t({\bm{v}})}T_{\bm{v}}\notin\mathcal{S}$, the commutation relation leads to $\langle S|T_{\bm{v}}|S\rangle=0$. Then, we generalize the Bell operator in Eq. (20) to any multipartite qudit stabilizer state. Consider the state $|S\rangle$ in $n$ qudits with stabilizer elements $S_{\bm{u}}=\omega^{[{\bm{a}},{\bm{u}}]}T_{\bm{u}}$ for all ${\bm{u}}\in\Sigma\subset\mathbb{Z}_{d}^{2n}$. For a unitary cube operator $U_{\nu}$ acting on the first qudit,
$\mathcal{B}_{S}=\sum_{{\bm{u}}\in\Sigma,t\in\mathbb{Z}_{d}}\chi^{\nu}_{{\bm{u}}}~{}~{}U^{\phantom{\dagger}}_{\nu}\,T_{u_{1}+(0,t)}\,U^{\dagger}_{\nu}\bigotimes_{i=2}^{n}T_{u_{i}}\,.$ (21)
@@
where $X_{(1/3)}=UXU^{\dagger}$, and $U=\mathrm{diag}(1,\omega^{2/3},\omega^{1/3})$. Note that these operators $O=X,X_{(1/3)}$ are not Hermitian but unitary yet represent valid measurements in terms of their Fourier transform with projectors $\Pi_{m}=d^{-1}\sum_{k\in\mathbb{Z}_{d}}\omega^{mk}O^{k}$ of outcome $m\in\mathbb{Z}_{d}$. The same holds for a deterministic value assignment. As a result, any lhv model attains $B_{3,\mathrm{lhv}}^{\max}=\max_{a_{0},a_{1},b_{0},b_{1}}(\omega^{a_{0}+b_{0}}+\omega^{2a_{0}+2b_{1}}+\omega^{2a_{1}+2b_{0}}+\omega^{a_{1}+b_{1}+1}+\,h.c.)=6+\omega+\omega^{2}=5$, where the maximum is taken over $a_{0},b_{0},a_{1},b_{1}\in\mathbb{Z}_{3}$, and is attained for $b_{1}=2$ and $a_{0},b_{0},a_{1}=0$. On the other hand, the quantum expectation value, $\mathrm{Tr}\left(\mathcal{B}_{3}\,|\Phi\rangle\langle\Phi|\right)=1+3(2\omega^{1/3}+\omega^{-2/3})+\,h.c.\,\approx 5.412\,,$ exceeds that of the classical model, $B_{3,\mathrm{lhv}}^{\max}<\mathrm{Tr}\left(\mathcal{B}_{3}\,|\Phi\rangle\langle\Phi|\right)$.
@@
We have constructed several families of qudit Bell inequalities for any odd prime dimension. First, we propose a novel form of general Bell inequalities where a bipartite entangled state exhibiting Wigner negativity acts as a witness for nonlocality. Similarly to the $\pi/8$-rotation leading to the CHSH inequality, our Bell operator consists of stabilizer elements under local rotation of a generalization of the qubit $\pi/8$-gate characterized in Howard2012 . We illustrate how the Bell inequalities can be regarded as a natural extension of the CHSH inequality. The constituting measurements test correlations related to the rotated states’ Wigner negativity, a witness of non-classicality. Our family of Bell operators, as well as its quantum and classical bound, is easy to analyze and can be adapted to any entangled stabilizer state, which also reduces the number of measurements. The measurements can be generalized to any non-Clifford unitary operator, and we specifically discuss those from higher-degree polynomials over finite fields, which we find to achieve larger Wigner negativity and thus increase the separation between quantum and classical models. Moreover, nonlocal correlations can emerge with qudit Pauli operators when rotated by unitary operators with a non-character spectrum in bipartite systems. These operators have an obvious advantage, which is resorting to phases $\omega^{q}$ that are beyond the description of any local value assignment that can only resort to characters $\omega^{k}$. Such operations might arise from interactions with higher-dimensional systems, but the feasibility of the implementation of any non-Clifford operator present in this work remains an open question. Its study heavily depends on the platform of experimental realization with possible candidates being super-conducting circuits, cold atoms and optical systems involving GKP (Gottesman–Kitaev–Preskill) states PhysRevA.64.012310 .